OpenStudy (anonymous):

Complex Rational Expressions: I will put the questions in the comment area! Some1 Please HELP me!

OpenStudy (anonymous):

3i-10 ----- 8i-5 11i^2 ----- -12i-8

OpenStudy (anonymous):

\[30\sqrt{-25}\times8\sqrt{-49}\] \[6\sqrt{-64}+12\sqrt{-36}\] \[4\sqrt{-13}-6\sqrt{-13}\]

OpenStudy (oaktree):

For the first one, you need to multiply by the conjugate, or 8i+5. So you get this:

OpenStudy (anonymous):

wouldn't it be 11i+5 ? Or is that wrong?

OpenStudy (oaktree):